# Interplay between Majorana and Shiba states in a minimal Kitaev chain coupled to a superconductor

### v1

@@ Appendix E Level Physics and Ultra-Strong coupling
Figure S 5: Non-local effects in SC transport as a function of the coupling to the SC lead. Upper (Lower) pannels, conductance for $\varepsilon_{2}/t=0$ ($\varepsilon_{2}/t=1$). Right column represents a close-up image signalling the transition to the ultra-strong coupling. The coupling to the normal lead is set to $\lambda_{n}/t=0.1$ in all pannels. Yellow contour marks the universal height value $G_{s}/G_{0}=(4-\pi)$.
@@ Appendix F Normal Transport
For the normal-superconducting case, it is convenient to define a different Gauge with time-independent tunnel couplings $\Sigma_{LR}$ and accounting for the chemical potential difference in the Keldysh GFs. As the coupling only involves one spin species, thus
$\begin{split}I_{NS}=\frac{e}{2h}\int\frac{1}{2}\textrm{tr}_{N}\big{\{}\sigma_{z}\Sigma_{LR}[G_{RR}^{-+}\Sigma_{RL}{\cal G}_{L}^{+-}\\ -\,G_{RR}^{+-}\Sigma_{RL}{\cal G}_{L}^{-+}]\big{\}}d\omega.\end{split}$ (24)
@@
$\check{G}=\begin{pmatrix}G_{LL}&&G_{LR}\\ G_{RL}&&G_{RR}\end{pmatrix}=\begin{pmatrix}{\cal G}_{L}^{-1}&&-\Sigma_{LR}\\ -\Sigma_{RL}&&{\cal G}_{R}^{-1}\end{pmatrix}^{-1}.$ (25)
From now on, we will use advanced GF and omit the explicit super-index. The non-equilibrium correlation functions can be expanded using the Langreth rules as
@@
${\cal G}^{+-}_{j}=F_{j}\big{(}{\cal G}_{j}-{\cal G}_{j}^{\dagger}\big{)}\ ,\quad{\cal G}^{+-}_{j}=(F_{j}-\mathbb{I})\big{(}{\cal G}_{j}-{\cal G}_{j}^{\dagger}\big{)},$ (26)
@@
$\sigma_{LR}$ $=$ $\sigma_{z}\Sigma_{LR}(\mathbb{I}+G_{LR}^{\dagger}\Sigma_{LR})({\cal G}_{R}-{\cal G}_{R}^{\dagger})(\mathbb{I}+\Sigma_{LR}^{\dagger}G_{LR})\Sigma_{LR}^{\dagger}({\cal G}_{L}-{\cal G}_{L}^{\dagger})(F_{R}-F_{L})$ (27)
@@
To disaggregate the different proccesses involved in transport, we explicitly expand in Nambu space the different terms of $\sigma_{LR}=\sigma_{LR}^{I}+\sigma_{LR}^{II}+\sigma_{LR}^{III}$ to compute the trace, where we have assumed that $\Sigma_{LR}=t_{e}\sigma_{z}$ and $t_{e}=\lambda_{n}$. Then, for example
$\sigma_{LR}^{I}=\sigma_{z}\Sigma_{LR}(\mathbb{I}+G_{LR}^{\dagger}\Sigma_{LR})A(\mathbb{I}+\Sigma_{LR}^{\dagger}G_{LR})\Sigma_{LR}^{\dagger}B(F_{R}-F_{L}),$ (28)
@@
$\sigma_{LR,ee}^{I}$ $=$ $t_{e}[(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{ee}-G^{\dagger}_{LR,he}t_{e}A_{he}](\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$ (29)
@@
where $A=({\cal G}_{R}-{\cal G}_{R}^{\dagger})$, and $B=({\cal G}_{L}-{\cal G}_{L}^{\dagger})$. From this term it is possible to extract the explicit expression for the different quasiparticle contributions to the electron current such
$\sigma_{Q_{1,ee}}={}$ $t_{e}(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{ee}(\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$
$+t_{e}G^{\dagger}_{LR,he}t_{e}A_{hh}t_{e}^{\dagger}G_{LR,he}t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)],$ (30a)
$\sigma_{Q_{2,ee}}={}$ $-t_{e}G^{\dagger}_{LR,he}t_{e}A_{he}(\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$
$-t_{e}(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{eh}t_{e}^{\dagger}G_{LR,he}t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)].$ (30b)
$\sigma_{A,ee}={}$ $t_{e}G^{\dagger}_{RR,he}t_{e}^{\dagger}B_{hh}t_{e}G_{RR,he}t_{e}^{\dagger}B_{ee}[n(\omega+eV)-n(\omega-eV)].$ (30c)
where the Andreev contribution to current is obtained from the remaining terms $\sigma_{A,ee}=\sigma_{LR,ee}^{II}+\sigma_{LR,ee}^{III}$. Equivalently, the hole contributions to the current can be computed as
$\sigma_{Q_{1,hh}}={}$ $-t_{e}(\mathbb{I}-G^{\dagger}_{LR,hh}t_{e})A_{hh}(\mathbb{I}-t_{e}^{\dagger}G_{LR,hh})t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]$
$-t_{e}G^{\dagger}_{LR,eh}t_{e}A_{ee}t_{e}^{\dagger}G_{LR,eh}t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)],$ (31a)
$\sigma_{Q_{2,hh}}={}$ $-t_{e}G^{\dagger}_{LR,eh}t_{e}A_{eh}(\mathbb{I}-t_{e}^{\dagger}G_{LR,hh})t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]$
$-t_{e}(\mathbb{I}-G^{\dagger}_{LR,hh}t_{e})A_{he}t_{e}^{\dagger}G_{LR,eh}t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)],$ (31b)
$\sigma_{A,hh}={}$ $t_{e}G^{\dagger}_{RR,eh}t_{e}^{\dagger}B_{ee}t_{e}G_{RR,eh}t_{e}^{\dagger}B_{hh}[n(\omega+eV)-n(\omega-eV)].$ (31c)
@@
$f^{\prime}_{L}=\frac{\partial}{\partial V}f(\omega\mp eV)\rightarrow\frac{\partial}{\partial V}(1-\Theta(\omega\mp eV))=\pm e\delta(\omega\mp eV),$ (32)
@@
$G_{NS}=-\frac{e^{2}}{2h}$ $\textrm{tr}\big{\{}\tilde{\sigma}_{A,ee}(-eV)+\tilde{\sigma}_{A,hh}(-eV)-\tilde{\sigma}_{Q_{1},hh}(-eV)-\tilde{\sigma}_{Q_{2},hh}(-eV)$ (33)
$+\tilde{\sigma}_{A,ee}(eV)+\tilde{\sigma}_{A,hh}(eV)+\tilde{\sigma}_{Q_{1},ee}(eV)+\tilde{\sigma}_{Q_{2},ee}(eV)\big{\}},$
where $\tilde{\sigma}_{\mu}$ represents the contributions to the current without the Fermi-Dirac terms, and the trace may take into account other possible orbital degrees of freedom.

### v3

@@ Appendix F Normal Transport
For the normal-superconducting case, it is convenient to define a different Gauge with time-independent tunnel couplings $\Sigma_{LR}$ and accounting for the chemical potential difference in the Keldysh GFs. As the coupling only involves one spin species, thus the current can be described in the $2\times 2$ Nambu basis
$\begin{split}I_{NS}=\frac{e}{2h}\int\frac{1}{2}\textrm{tr}_{N}\big{\{}\sigma_{z}\Sigma_{LR}[G_{RR}^{-+}\Sigma_{RL}{\cal G}_{L}^{+-}\\ -\,G_{RR}^{+-}\Sigma_{RL}{\cal G}_{L}^{-+}]\big{\}}d\omega\,.\end{split}$ (S 23)
@@
$\check{G}=\begin{pmatrix}G_{LL}&&G_{LR}\\ G_{RL}&&G_{RR}\end{pmatrix}=\begin{pmatrix}{\cal G}_{L}^{-1}&&-\Sigma_{LR}\\ -\Sigma_{RL}&&{\cal G}_{R}^{-1}\end{pmatrix}^{-1}.$ (S 24)
From now on, we will use advanced GF and omit the explicit super-index. The non-equilibrium correlation functions can be expanded using the Langreth rules [77, 78, 79] as
@@
${\cal G}^{+-}_{j}=F_{j}\big{(}{\cal G}_{j}-{\cal G}_{j}^{\dagger}\big{)}\ ,\quad{\cal G}^{+-}_{j}=(F_{j}-\mathbb{I})\big{(}{\cal G}_{j}-{\cal G}_{j}^{\dagger}\big{)}\,,$ (S 25)
@@
$\kappa_{LR}$ $=$ $\sigma_{z}\Sigma_{LR}(\mathbb{I}+G_{LR}^{\dagger}\Sigma_{LR})({\cal G}_{R}-{\cal G}_{R}^{\dagger})(\mathbb{I}+\Sigma_{LR}^{\dagger}G_{LR})\Sigma_{LR}^{\dagger}({\cal G}_{L}-{\cal G}_{L}^{\dagger})(F_{R}-F_{L})$ (S 26)
@@
To disaggregate the different processes involved in transport, we explicitly expand in Nambu space the different terms of $\kappa_{LR}=\kappa_{LR}^{I}+\kappa_{LR}^{II}+\kappa_{LR}^{III}$ to compute the trace, where we have assumed that $\Sigma_{LR}=t_{e}\sigma_{z}$ and $t_{e}=\lambda_{n}$. Then, for example
$\kappa_{LR}^{I}=\sigma_{z}\Sigma_{LR}(\mathbb{I}+G_{LR}^{\dagger}\Sigma_{LR})A(\mathbb{I}+\Sigma_{LR}^{\dagger}G_{LR})\Sigma_{LR}^{\dagger}B(F_{R}-F_{L})\,,$ (S 27)
@@
$\kappa_{LR,ee}^{I}$ $=$ $t_{e}[(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{ee}-G^{\dagger}_{LR,he}t_{e}A_{he}](\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$ (S 28)
@@
where $A=({\cal G}_{R}-{\cal G}_{R}^{\dagger})$, and $B=({\cal G}_{L}-{\cal G}_{L}^{\dagger})$. From this term it is possible to extract the explicit expression for the different quasiparticle contributions to the electron current, such
$\kappa_{Q_{1,ee}}={}$ $t_{e}(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{ee}(\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$
$+t_{e}G^{\dagger}_{LR,he}t_{e}A_{hh}t_{e}^{\dagger}G_{LR,he}t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]\,,$ (S 29a)
$\kappa_{Q_{2,ee}}={}$ $-t_{e}G^{\dagger}_{LR,he}t_{e}A_{he}(\mathbb{I}+t_{e}^{\dagger}G_{LR,ee})t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]$
$-t_{e}(\mathbb{I}+G^{\dagger}_{LR,ee}t_{e})A_{eh}t_{e}^{\dagger}G_{LR,he}t_{e}^{\dagger}B_{ee}[n(\omega)-n(\omega-eV)]\,.$ (S 29b)
$\kappa_{A,ee}={}$ $t_{e}G^{\dagger}_{RR,he}t_{e}^{\dagger}B_{hh}t_{e}G_{RR,he}t_{e}^{\dagger}B_{ee}[n(\omega+eV)-n(\omega-eV)]\,.$ (S 29c)
where the Andreev contribution to current is obtained from the remaining terms $\kappa_{A,ee}=\kappa_{LR,ee}^{II}+\kappa_{LR,ee}^{III}$. Equivalently, the hole contributions to the current can be computed as
$\kappa_{Q_{1,hh}}={}$ $-t_{e}(\mathbb{I}-G^{\dagger}_{LR,hh}t_{e})A_{hh}(\mathbb{I}-t_{e}^{\dagger}G_{LR,hh})t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]$
$-t_{e}G^{\dagger}_{LR,eh}t_{e}A_{ee}t_{e}^{\dagger}G_{LR,eh}t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]\,,$ (S 30a)
$\kappa_{Q_{2,hh}}={}$ $-t_{e}G^{\dagger}_{LR,eh}t_{e}A_{eh}(\mathbb{I}-t_{e}^{\dagger}G_{LR,hh})t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]$
$-t_{e}(\mathbb{I}-G^{\dagger}_{LR,hh}t_{e})A_{he}t_{e}^{\dagger}G_{LR,eh}t_{e}^{\dagger}B_{hh}[n(\omega)-n(\omega-eV)]\,,$ (S 30b)
$\kappa_{A,hh}={}$ $t_{e}G^{\dagger}_{RR,eh}t_{e}^{\dagger}B_{ee}t_{e}G_{RR,eh}t_{e}^{\dagger}B_{hh}[n(\omega+eV)-n(\omega-eV)]\,.$ (S 30c)
@@
$f^{\prime}_{L}=\frac{\partial}{\partial V}f(\omega\mp eV)\rightarrow\frac{\partial}{\partial V}(1-\Theta(\omega\mp eV))=\pm e\delta(\omega\mp eV)\,,$ (S 31)
@@
$G_{NS}=-\frac{e^{2}}{2h}$ $\textrm{tr}\big{\{}\tilde{\kappa}_{A,ee}(-eV)+\tilde{\kappa}_{A,hh}(-eV)-\tilde{\kappa}_{Q_{1},hh}(-eV)-\tilde{\kappa}_{Q_{2},hh}(-eV)$ (S 32)
$+\tilde{\kappa}_{A,ee}(eV)+\tilde{\kappa}_{A,hh}(eV)+\tilde{\kappa}_{Q_{1},ee}(eV)+\tilde{\kappa}_{Q_{2},ee}(eV)\big{\}}\,,$
where $\tilde{\kappa}_{\mu}$ represents the contributions to the current without the Fermi-Dirac terms, and the trace may take into account other possible orbital degrees of freedom.
Appendix G Ultra-Strong Coupling Regime
Figure S 6 illustrates various normal conductance profiles for different values of $\Gamma_{s}$ considering the interplay between MZM and YSR states when the system deviates from the sweet spot. In all cases we consistently see that all subgap states tend towards zero energy for large enough couplings to the SC lead. Therefore, the gap to excited states tends towards zero with the coupling to the superconducting lead.
As discussed in the main text, the normal conductance at the sweet spot in Fig. S 6(a) shows a quantized YSR states coexisting with a quantized zero-bias peak. Remarkably, in the ultra-strong coupling regime the quantized zero-bias conductance peak persist, despite the overlap of YSR and Majorana states.
Figure S 6(b,c) shows the YSR-MZM physics when one of the QDs in the chain is detuned. In the case of $\varepsilon_{2}\neq 0$, shown in FIG. S 6(b), it should be noticed the absence of quantized transport response for the YSR states, resulting in a vanishing normal transport signal for larger couplings. Nonetheless, the quantized zero-bias peak remains intact, indicating the robustness of the MZM peak against this detuning. Conversely, FIG. S 6(c), considers the case where $\varepsilon_{1}\neq 0$. Here, a distinct quantized peak associated with the YSR states emerges, accompanied by a zero-bias dip. As the coupling to the SC lead is further increased, the low-bias conductance exhibits an increase as the YSR energy decreases. When both subgap states merge, the quantized values area reached for both broadened peaks, while the zero-bias dip is preserved.
Lastly, Fig. S 6(d) addresses the $t\neq\Delta$ case, revealing the MZM as a non-zero energy subgap state that displays quantized conductance. In the strong coupling, the YSR state begins to detach from the continuum. This non-trivial subgap structure aligns with the evolution of poles detailed in App. A. When the subgap states are well isolated between them and the continuum, they show quantized conductance behaviour. However, at even higher values of $\Gamma_{s}$, these states merge as they converge toward zero-bias. This merging effect results in a pronounced zero-bias dip and a decrease in the overall transport signal as the coupling to the SC lead continues to increase, illustrating the complex interplay between both subgap states.
Figure S 6: Interplay between YSRs and MZMs in normal conductance as a function of the coupling to the SC lead. (a) Conductance cuts in the sweet spot configuration $t/\Delta=1$ and $\varepsilon_{1}=\varepsilon_{2}=0$. The rest of the panels shows situations out of the sweet spot such (b) $\varepsilon_{2}\neq 0$ (c) $\varepsilon_{1}\neq 0$ and (d) $t/\Delta=0.5$. Inset shows a close-up image around zero-bias. The coupling to the normal lead is set to $\lambda_{n}/t=1$ in all panels.